OpenStudy (anonymous):

correction, the problem is 2x+4y-1=0

OpenStudy (anonymous):

rearrange to get slope-intercept form 4y = -2x + 1 y = (-1/2)x + 1/4 slope = -1/2 and y intercept is 1/4 ( where x = 0)

OpenStudy (anonymous):

So, why does -2x become -1/2x? I'm sorry but I'm really bad at math.

OpenStudy (anonymous):

ok - i have divided each term by 4 to get y only on the left so i need to do same to -2x and the 1 -2/4 = -1/2 and 1/4 slope intercept fore is y = mx + c where m = slope and c = y intercept

OpenStudy (anonymous):

and 1 over 4 = 1/4

OpenStudy (anonymous):

Oh, ok, I got it. Thanks!!!

OpenStudy (anonymous):

no probs